OpenStudy (anonymous):

I have no idea how to do this. Please help me? It's multiple choice. Simplify completely (8x-16)/(x^2-13+22) and find the restrictions on the variable

OpenStudy (anonymous):

\[\frac{ 8x-16 }{ x^2-13+22 }\]

OpenStudy (anonymous):

\[\frac{ 8x-16 }{ x^{2}+9 } ?\]

OpenStudy (anonymous):

\[A. \frac{ x-2 }{ ?x-11}, x \neq11\]

OpenStudy (anonymous):

Restrictions: \[x^{2} + 9 \neq 0\]

OpenStudy (anonymous):

\[B. \frac{ x-2 }{ x-11 }, x \neq2, x \neq11\]

OpenStudy (anonymous):

\[C. \frac{ 8 }{ x-11 }, x \neq11\]

OpenStudy (anonymous):

\[D. \frac{ 8 }{ x-11}, x \neq2, x \neq11\]

OpenStudy (anonymous):

@Soumyadeep That isn't an option... :(

OpenStudy (anonymous):

Option is D

OpenStudy (anonymous):

\[\frac{ 8(x-2) }{ (x-2)(x-11) }\]

OpenStudy (anonymous):

Thank you so much! :D